Assume that adults have IQ scores that are normally distributed with a mean of

95.4

and a standard deviation of

25

Find the probability that a randomly selected adult has an IQ greater than

142.7

​(Hint: Draw a​ graph.)
Answer #1

Solution :

Given that,

mean = = 95.4

standard deviation = =25

P (x > 142.7 )

= 1 - P (x < 142.7 )

= 1 - P ( x -  / ) < ( 142.7 - 95.4 / 25)

= 1 - P ( z < 47.3 / 25 )

= 1 - P ( z < 1.89 )

Using z table

= 1 - 0.9706

= 0.0294

Probability = 0.0294